OpenStudy (anonymous):

3m(-6m – 2m2p – 10p)

Directrix (directrix):

Distribute the (3m) over the quantity (-6m – 2m^2p – 10p). 3m(-6m – 2m^2p – 10p) = (3m) * (-6m) + 3m*(– 2m^2p) + 3m * (-10p) = ? Starting here, multiply these terms and post your answer. Please leave your work as the sum of the three terms you get after multiplying. After we check that, we'll see if any of those three terms are like terms. (3m) * (-6m) + 3m*(– 2m^2p) + 3m * (-10p) = ?